OpenStudy (anonymous):

Differentiate ln((4-x)/(4+x))

OpenStudy (cj7529):

first separate the logs ln[(4-x)/(4+x)]= =ln(4-x) - ln(4+x) then differentiate the two logs individually =-1/(4-x) - 1/(4+x) and you can then combine the fraction but its not necessary.

OpenStudy (anonymous):

okay, thanks!